# Keep the beat going: Automatic drum transcription with momentum

### v1

@@
Abstract
A simple, interpretable way to perform automatic drum transcription is by factoring the magnitude spectrogram of a recorded musical piece using a partially fixed nonnegative matrix factorization. There are two natural ways to optimize the nonnegative matrix factorization, including a multiplicative update rule and projected gradient descent with momentum. The methods differ in their empirical accuracies and theoretical convergence guarantees. This paper summarizes the methods and their time complexities, and it applies the methods to the ENST-Drums data set and an original recording from the author’s band, evaluating the empirical accuracy with respect to ground-truth drum annotations. The results indicate that projected gradient descent with momentum leads to higher accuracy for a fixed runtime, and it satisfies stronger convergence guarantees.
1 Introduction
Nonnegative matrix factorization (NMF) is a data science approach that approximates a matrix as a product of two nonnegative matrices. NMF was introduced in 1994 by Paatero and Tapper as “positive matrix factorization” [14], and it was popularized by Lee and Seung [9] in 2000. NMF has many applications in music analysis [4], image representation [12], recommender systems [13], and genomics [18].
This work will focus on applying NMF to automatic drum transcription, a subfield within automatic music transcription [16]. An automatic drum transcription system could be useful to drummers as a tool for notating their grooves, as well as learning beats from recorded musical pieces. Previous work [15, 20, 19] has applied NMF to drum source detection, separation, and transcription. Researchers have also extended the NMF approach to automatic drum transcription using support vector machines [7] and convolutional neural networks [8].
This paper will investigate the partially fixed NMF method for automatic drum transcription developed in [20, 19]. There are two natural strategies to optimize the partially fixed NMF: (1) a multiplicative update rule based on gradient descent and (2) a projected gradient descent approach with Nesterov momentum. The main goal of this work is to compare the convergence of the two optimization algorithms. The results show that projected gradient descent with Nesterov momentum has favorable convergence guarantees and outperforms the multiplicative update rule in its empirical accuracy with a fixed runtime.
The paper is organized as follows. Section 2 describes the problem formulation, Section 3 summarizes optimization methods for NMF, Section 4 applies the methods to automatic drum transcription of the ENST-Drums database [5] and a demo of “Every Now and Then” by the author’s band, Threadbare [17], and Section 5 offers concluding remarks.
1.1 Notation
Scalars are in regular typeface: $n,d,L$. Vectors are in bold lower-case letters: $\bm{x},\bm{y}$. Matrices are in bold capital letters: $\bm{M},\bm{N}$. The $i$th entry of a vector $\bm{x}$ is $\bm{x}(i)$. The $(i,j)$ entry of a matrix $\bm{M}$ is $\bm{M}(i,j)$, the $i$th row is $\bm{M}(i,\cdot)$, and the $j$th column is $\bm{M}(\cdot,j)$. The transpose is $\bm{M}^{\top}$, the Moore-Penrose pseudoinverse is $\bm{M}^{+}$, the spectral norm is $\lVert\bm{M}\rVert$, and the Frobenius norm is $\lVert\bm{M}\rVert_{\rm F}$ or $\lVert\bm{x}\rVert_{2}$. The space of nonnegative-valued $m\times n$ matrices is denoted +$\mathbb{R}^{m\times n}_{+}$. The element-wise product of two vectors $\bm{x},\bm{y}$ is $\bm{x}\odot\bm{y}$. The element-wise product of two matrices $\bm{M},\bm{N}$ is $\bm{M}\odot\bm{N}$, and the element-wise division is $\frac{\bm{M}}{\bm{N}}$.
2 Problem formulation
The classic NMF problem is
(1) $\min_{\bm{W},\bm{H}}\lVert\bm{V}-\bm{WH}\rVert_{\rm F}^{2}$
where $\bm{V}$ is an $m\times n$ matrix and $\bm{W}\in\mathbb{R}^{m\times r}$ and $\bm{H}\in\mathbb{R}^{r\times n}$ are constrained to have all nonnegative entries.
@@ 2 Problem formulation
In the context of automatic music transcription, $\bm{V}$ is the magnitude spectrogram matrix of a piece of music that has $m$ frequency bins and $n$ time frames. $\bm{W}$ is an $m\times r$ dictionary matrix that contains a basis of frequencies for the song, and $\bm{H}$ is an $r\times n$ activation matrix containing the temporal onsets of each component of $\bm{W}$ [16]. In this context, it is natural that $\bm{V}$ is approximated by the product of $\bm{W}$ and $\bm{H}$ matrices that both have nonnegative entries.
In 2010, Yoo et al. [20] introduced a method to separate drums from the rest of a musical track by assigning certain columns of the dictionary matrix and certain rows of the activation matrix to the drums. In 2015, Wu and Lerch [19] applied a similar approach, where they defined dictionary and activation matrices
@@
Here $\bm{W}_{D}$ is an $m\times r_{D}$ matrix whose columns indicate the frequency bases of each part of the drum kit. Typically the rank $r_{D}$ is the number of drum parts (snare drum, bass drum, hi-hat, etc.) to be transcribed. $\bm{W}_{H}$ contains the frequency bases of the harmonic instruments in the song, and its rank, $r_{H}$, is tuned to balance accuracy with interpretability. $\bm{H}_{D}$ is an $r_{D}\times n$ matrix that indicates when the components of $\bm{W}_{D}$ are active. Similarly, $\bm{H}_{H}$ is the $r_{H}\times n$ activation matrix for the harmonic instruments of the song.
In their “partially fixed NMF” approach, Wu & Lerch [19] train the matrix $\bm{W}_{D}$ using the ENST drums data set [5], and they keep this matrix fixed as $\bm{H}_{D},\bm{W}_{H},\bm{H}_{H}$ are updated. Therefore, the problem of interest Eq. 1 is modified and reformulated as minimizing a loss function
(2) $F(\bm{H}_{D},\bm{W}_{H},\bm{H}_{H})=\tfrac{1}{2}\lVert\bm{V}-(\bm{W}_{D}\bm{H}_{D}+\bm{W}_{H}\bm{H}_{H})\rVert_{\rm F}^{2},$
subject to the constraints that $\bm{H}_{D}$, $\bm{W}_{H}$, $\bm{H}_{H}$ are all nonnegative-valued. Yoo et al. [20] also consider methods of adapting $\bm{W}_{D}$ as the algorithm iterates, but this study will only consider fixed $\bm{W}_{D}$.
The major benefit of the partially fixed NMF approach is that having a pre-trained $\bm{W}_{D}$ allows for greater interpretability of the components. For example, the first column might represent the snare drum, the second column might represent the bass drum, etc. This ordering is necessary for transforming the data into human-readable drum notation. In practice, a drummer could calibrate $\bm{W}_{D}$ in the automatic drum transcription system by recording some drum hits to create a frequency basis for each part of a drum kit.
3 Summary of optimization methods
It is extremely challenging to find the global optimizer of the partially fixed NMF problem Eq. 2, since the loss function is nonconvex. Nevertheless, researchers have proposed finding a local optimizer using an alternating minimization [20]:
(3) $\bm{H}_{D}^{t+1}$ $=\operatorname*{arg\,min}_{\bm{H}_{D}\geq 0}F(\bm{H}_{D},\bm{W}_{H}^{t},\bm{H}_{H}^{t})$
$\bm{W}_{H}^{t+1}$ $=\operatorname*{arg\,min}_{\bm{W}_{H}\geq 0}F(\bm{H}_{D}^{t+1},\bm{W}_{H},\bm{H}_{H}^{t})$
$\bm{H}_{H}^{t+1}$ $=\operatorname*{arg\,min}_{\bm{H}_{H}\geq 0}F(\bm{H}_{D}^{t+1},\bm{W}_{H}^{t+1},\bm{H}_{H})$
Each subproblem in Eq. 3 is convex, so it can be solved relatively efficiently. There are two natural strategies to solve each subproblem based on a multiplicative update rule (MUR) or projected gradient descent with Nesterov momentum (NeNMF). The rest of this section will derive the MUR and NeNMF methods in detail and derive the best available mathematical convergence guarantees, building on techniques in [9, 6].
3.1 Multiplicative update rule
Lee and Seung [9] studied the classic NMF problem Eq. 1, and they proved the following result.
Theorem 3.1 (Convergence of original MUR algorithm [9]).
@@ 3.1 Multiplicative update rule
The square Frobenius norm $\|\bm{V}-\bm{WH}\|^{2}$ is nonincreasing under the multiplicative update rule
(4) $\bm{H}\leftarrow\bm{H}\odot\frac{\bm{W}^{\top}\bm{V}}{\bm{W}^{\top}\bm{WH}},\qquad\bm{W}\leftarrow\bm{W}\odot\frac{\bm{VH}^{\top}}{\bm{WHH}^{\top}}.$
@@
In a critique of the MUR method, Lin [11] pointed out that a denominator term might equal zero, in which case the method is ill-defined. However, a zero denominator only arises in pathological settings, and Theorem 3.1 remains valid as long as the denominators stay positive.
Algorithm 1 MUR for partially fixed NMF
3:  for $t=1,...,T$ do
7:  end for
8:  $\bm{H}_{D}=\bm{H}_{D}^{T}$, $\bm{W}_{H}=\bm{W}_{H}^{T}$, $\bm{H}_{H}=\bm{H}_{H}^{T}$.
9:  return $\bm{H}_{D},\bm{W}_{H},\bm{H}_{H}$
In 2010, Yoo et al. [20] proposed solving the partially fixed NMF problem Eq. 2 using a variant of the multiplicative update rule Eq. 4. See Algorithm 1 for pseudocode derived from their work. As a limitation, however, the paper [20] did not provide a complete convergence analysis. To fill in this gap, the current paper proves a theoretical guarantee for MUR method by elaborating on the proof strategy in [9, Thm. 1].
Theorem 3.2 (Convergence of MUR algorithm for partially fixed NMF).

### v2

@@
Abstract.
How can we process a piece of recorded music to detect and visualize the onset of each instrument? A simple, interpretable approach is based on partially fixed nonnegative matrix factorization (NMF). Yet despite the method’s simplicity, partially fixed NMF is challenging to apply because the associated optimization problem is high-dimensional and non-convex. This paper explores two optimization approaches that preserve the nonnegative structure, including a multiplicative update rule and projected gradient descent with momentum. These techniques are derived from the previous literature, but they have not been fully developed for partially fixed NMF before now. Results indicate that projected gradient descent with momentum leads to the higher accuracy among the two methods, and it satisfies stronger local convergence guarantees.
Key words and phrases:
Nonnegative matrix factorization, Nesterov momentum, automatic drum transcription
1991 Mathematics Subject Classification:
65F55
1. Introduction
Drums are the rhythmic backbone of a song, contributing to our perception and enjoyment of music. A typical drum kit consists of many parts— such as a snare drum, bass drum, and multiple cymbals— which are played simultaneously to create a drum groove. This paper considers the computational inverse problem: given a recorded musical piece, how can we extract the drum signals that created it?
Automatic drum transcription (ADT) is the act of extracting and notating drum grooves from recorded music. In ADT, the data are transformed into spectrograms, time-frequency representations calculated with the short-time Fourier transform. The spectrograms are analyzed via ADT systems to precisely detect the timing of the drum hits. Yet a challenge for ADT systems is that drums are inharmonic instruments, meaning they do not produce pure musical tones like a piano or guitar. Moreover, each drum or cymbal sounds different based on physical aspects such as its size, material, and tension. This makes drum transcription more difficult than the transcription of a piano, for example, where each A4 note can be fully characterized by its fundamental frequency of 440Hz. A typical ADT system needs to be adapted to the individual drum kit.
Ultimately, an ADT system could be useful to drummers as a tool for notating their grooves or learning beats from recorded musical pieces. The ideal ADT system would be fully interpretable, identifying the frequency representations of each part of the drum kit. This work constructs an interpretable ADT method based on partially fixed nonnegative matrix factorization (NMF), and it identifies an efficient numerical method for the partially fixed NMF optimization. See fig. 1 for an illustration of the interpretable ADT system constructed in this work. The optimization approach is based on established methods [14, 10], but the theory and implementation of these methods need to be modified for the partially fixed NMF problem. The rest of this introductory section describes the partially fixed NMF problem and the optimization approach in more detail.
Figure 1. Automatic drum transcription (ADT) system applied to the first 30 seconds of the song “Every Now and Then” to visualize the onsets of snare and bass drums.
1.1. Constructing an interpretable ADT model
To give a brief background, NMF approximates a nonnegative-valued data matrix $\bm{V}$ as the product of two nonnegative-valued matrices, called a dictionary matrix $\bm{W}$ and an activation matrix $\bm{H}$.
$\bm{V}\approx\bm{W}\bm{H}.$
@@ 1.1. Constructing an interpretable ADT model
In the context of automatic music transcription, $\bm{V}$ is the $m\times n$ magnitude spectrogram matrix that has $m$ frequency bins and $n$ time frames. $\bm{W}$ is an $m\times r$ dictionary matrix that contains a basis of frequencies for the song, and $\bm{H}$ is an $r\times n$ activation matrix that contains the temporal onsets of each frequency component. In this context, it is natural that $\bm{V}$ is approximated by the product of $\bm{W}$ and $\bm{H}$ matrices that both have nonnegative entries.
Paatero and Tapper advocated for NMF in 1994 and called it “positive matrix factorization” [21]. Later, Lee and Seung popularized the method in 2000 [14]. NMF has many applications in music analysis [6], image representation [19], recommender systems [20], and genomics [25]. Section 2 reviews existing works that apply NMF to ADT.
As the main limitation, the conventional NMF model is underspecified. Multiple columns of $\bm{W}$ can represent the same instrument or multiple instruments can be represented in the same column. Therefore, our work will investigate a more highly specified model called partially fixed NMF [28, 26], where the dictionary and activation matrices are factored as
@@
Here $\bm{W}_{D}$ is an $m\times r_{D}$ matrix, where the rank $r_{D}$ is the number of drum parts to be transcribed, and each column contains one drum component that we want to learn, e.g., snare drum, bass drum, and hi-hat cymbal. In practice, a drummer could calibrate $\bm{W}_{D}$ by recording some drum hits to create a frequency basis for each part of a drum kit. For example, fig. 2 illustrates the columns of $\bm{W}_{D}$ corresponding to snare and bass drums in green and red lines. This set of columns stays fixed for the duration of the optimization process. $\bm{H}_{D}$ is an $r_{D}\times n$ matrix that contains the temporal onsets of the known drum instruments needed for transcription. For example, fig. 1 illustrates the rows of $\bm{H}_{D}$ corresponding to snare and bass drums. Last, $\bm{W}_{H}$ contains the frequencies of the harmonic instruments in the song, and the rank $r_{H}$ is tuned to balance accuracy with interpretability. Similarly, $\bm{H}_{H}$ is the $r_{H}\times n$ activation matrix for the harmonic instruments of the song. $\bm{W}_{H}\bm{H}_{H}$ essentially absorbs all non-drum noise from the spectrogram.
Figure 2. Plot of the magnitudes of the frequencies for the basis components in $\bm{W}_{D}$ for the snare and bass drums for “Every Now and Then.”
1.2. Addressing the optimization challenge
It is challenging to optimize the objective function in the partially fixed NMF model, which is given by
$F(\bm{H}_{D},\bm{W}_{H},\bm{H}_{H})=\frac{1}{2}\lVert\bm{V}-(\bm{W}_{D}\bm{H}_{D}+\bm{W}_{H}\bm{H}_{H})\rVert_{\rm F}^{2}.$
The objective function is high-dimensional and non-convex. Nonetheless, the NMF community has identified an optimization approach using alternating minimization that works well in practice. Freeze two matrices in the factorization and optimize the other matrix, which becomes a convex optimization problem. Then trade the roles of the matrices. Repeat as many times as desired.
$\bm{H}_{D}^{t+1}$ $=\operatornamewithlimits{argmin}_{\bm{H}_{D}\geq 0}F(\bm{H}_{D},\bm{W}_{H}^{t},\bm{H}_{H}^{t})$
$\bm{W}_{H}^{t+1}$ $=\operatornamewithlimits{argmin}_{\bm{W}_{H}\geq 0}F(\bm{H}_{D}^{t+1},\bm{W}_{H},\bm{H}_{H}^{t})$
$\bm{H}_{H}^{t+1}$ $=\operatornamewithlimits{argmin}_{\bm{H}_{H}\geq 0}F(\bm{H}_{D}^{t+1},\bm{W}_{H}^{t+1},\bm{H}_{H}).$
There is no guarantee the approach will converge to a global minimizer, but the square Frobenius norm error can be made to decrease at each step.
The optimization part of this paper focuses on a subproblem in the alternating minimization: how do we optimize one matrix when the others are fixed? There are at least two natural strategies that preserve the nonnegative structure: (1) a multiplicative update rule based on gradient descent [14] and (2) a projected gradient descent approach with Nesterov momentum [10]. We need to execute these methods carefully to make sure the iterates remain nonnegative, and we need to carefully summarize the relevant theory to support each approach. This work will adapt the two optimization strategies to the problem of partially fixed NMF, developing theory and algorithms. The results show that projected gradient descent with Nesterov momentum has favorable convergence guarantees and outperforms the multiplicative update rule in its empirical accuracy with a fixed runtime.
1.3. Outline
The paper is organized as follows. Section 2 reviews existing work, section 3 describes the partially fixed NMF problem and formulates optimization methods, section 4 applies the methods to automatic drum transcription of the ENST-Drums database [7] and a demo of “Every Now and Then” by the author’s band, Threadbare [24], and section 5 concludes.
1.4. Notation
Scalars are in regular typeface: $n,d,L$. Vectors are in bold lower-case letters: $\bm{x},\bm{y}$. Matrices are in bold capital letters: $\bm{M},\,\bm{N}$. The $i$th entry of a vector $\bm{x}$ is $\bm{x}(i)$. The $(i,j)$ entry of a matrix $\bm{M}$ is $\bm{M}(i,j)$, the $i$th row is $\bm{M}(i,\cdot)$, and the $j$th column is $\bm{M}(\cdot,j)$. The transpose is $\bm{M}^{\top}$, the Moore-Penrose pseudoinverse is $\bm{M}^{+}$, the spectral norm is $\lVert\bm{M}\rVert$, and the Frobenius norm is $\lVert\bm{M}\rVert_{\rm F}$. The space of nonnegative-valued $m\times n$ matrices is denoted $\mathbb{R}^{m\times n}_{+}$. The element-wise product of two vectors $\bm{x},\bm{y}$ is $\bm{x}\odot\bm{y}$. Similarly, the element-wise product of two matrices $\bm{M},\,\bm{N}$ is $\bm{M}\odot\bm{N}$, and the element-wise division is $\frac{\bm{M}}{\bm{N}}$.
2. Literature Review
This section reviews ADT models from the literature. Drawbacks to some existing models are the large amount of data required for optimization, the slow optimization speed, and the lack of human interpretability.
2.1. A variety of ADT models
ADT has been performed with several neural network-based models. For example, Jacques and Roebel [12] trained three convolutional neural networks using a data set of over 400 musical pieces — one each for detecting the hi-hat, snare drum, and bass drum — and achieved a high degree of accuracy. As a drawback, however, neural networks are not the most interpretable model choice. Therefore, many researchers have applied NMF for more interpretable transcriptions of drum music [22, 28, 26] or polyphonic music [23, 27, 5]. NMF takes advantage of the inherent nonnegativity and low-rank properties of magnitude spectrograms, and a possible advantage of NMF models is that they require little training data.
One popular version of the NMF model, called “convolutive NMF” (CNMF), uses a convolution operator to account for local dependencies between nearby time frames. Smaragdis [22] used the CNMF model to transcribe monophonic drum data. Later, H. Wu et al. [27] used CNMF to transcribe piano music. The CNMF method is not perfect, however. A drawback noted by H. Wu et al. is the long computation time needed for transcriptions, although recent work has pursued more robust optimization algorithms for CNMF [4].
Any NMF model can be used with a variety of loss functions. The Frobenius norm loss function is common [14, 10, 18], and two other popular loss functions in audio source separation are the Kullback-Leibler divergence [15, 14] and Itakura-Saito divergence [15, 6]. The latter two loss functions are logarithmically scaled, mirroring human audio perception [15]. Here we use the simplest, most common Frobenius norm loss function, but exploring alternative loss functions is an interesting topic for future work.
2.2. Partially fixed NMF
As a limitation, traditional NMF models are underspecified, so they might assign multiple columns of the dictionary matrix to the same instrument or assign one column to multiple instruments. In order to transcribe sheet music in this context, a drummer would need to train an auxiliary model identifying which instruments produce which onsets. This work focuses on a more highly specified NMF model called “partially fixed NMF” that was proposed by Yoo et al. [28] and later extended by C.-W. Wu and Lerch [26]. The major benefit of the partially fixed NMF is interpretability. In partially fixed NMF, each drum instrument is assigned to a single column of the dictionary matrix, and the drum onsets are described by the corresponding rows of the activation matrix. For example, the first column of the dictionary matrix might represent the snare drum, the second column might represent the bass drum, etc. and then the first row of the activation matrix would contain the snare drum onsets and the second row would contain the bass drum onsets. This ordering is helpful for transforming the data into human-readable drum notation.
2.3. Optimizers for NMF
There are many optimizers for NMF problems, the most classic and prevalent being the multiplicative update rule (MUR) presented by Lee and Seung [14]. Projected gradient descent is another first-order method suitable for high-dimensional optimization [18], and Guan et al. have developed a projected gradient descent method with Nesterov momentum for NMF called “NeNMF” [10]. Other NMF optimizers include alternating least squares [21], alternating nonnegative least squares [13], the alternating direction method of multipliers [1], hierarchical alternating least squares [9], and proximal block coordinate descent [11]. Here we focus on analyzing, applying, and comparing MUR and NeNMF. As motivation, MUR is the most classic and widely used optimizer, while NeNMF converges more rapidly than MUR in the experiments of [10] and [8, Sec. 8.4]. MUR and NeNMF are highly flexible NMF optimization methods that can be adapted to a range of optimization problems, as emphasized in [8, Tab. 8.2]. Therefore, it makes sense to develop these methods first when optimizing the partially fixed NMF. Nonetheless, comparisons between more optimizers would be interesting further work.
3. Methods and optimization techniques
This section will describe the MUR and NeNMF optimization methods in detail and derive the best available mathematical convergence guarantees, building on techniques in [14, 10].
3.1. Multiplicative update rule
Lee and Seung [14] studied the classic NMF problem
$\min_{\bm{W},\bm{H}}\frac{1}{2}\lVert\bm{V}-\bm{WH}\rVert_{\rm F}^{2},$ (1)
and they proved the following result.
Theorem 1 (Convergence of original MUR algorithm [14]).
@@ 3.1. Multiplicative update rule
The square Frobenius norm $\lVert\bm{V}-\bm{WH}\rVert_{\rm F}^{2}$ is nonincreasing under the multiplicative update rule
$\bm{H}\leftarrow\bm{H}\odot\frac{\bm{W}^{\top}\bm{V}}{\bm{W}^{\top}\bm{WH}},\qquad\bm{W}\leftarrow\bm{W}\odot\frac{\bm{VH}^{\top}}{\bm{WHH}^{\top}}.$ (2)
@@
In a critique of the MUR method, Lin [17] pointed out that terms in the denominator might equal zero, in which case the method is ill-defined. However, zero denominator terms only arise in pathological settings, and 1 remains valid as long as the denominators stay positive.
In 2010, Yoo et al. [28] proposed solving the partially fixed NMF problem
$\min_{\bm{H}_{D},\bm{W}_{H},\bm{H}_{H}}\frac{1}{2}\lVert\bm{V}-(\bm{W}_{D}\bm{H}_{D}+\bm{W}_{H}\bm{H}_{H})\rVert_{\rm F}^{2}.$ (3)
using a variant of the multiplicative update rule eq. 2. Their approach is a little subtle, since we cannot simply replace $\bm{W}$ with $\bm{W}_{D}$, $\bm{H}$ with $\bm{H}_{D}$, and $\bm{V}$ with $\bm{V}-\bm{W}_{H}\bm{H}_{H}$ in eq. 2 to produce an update
$\bm{H}_{D}\leftarrow\bm{H}_{D}\odot\frac{\bm{W}_{D}^{\top}[\bm{V}-\bm{W}_{H}\bm{H}_{H}]}{\bm{W}_{H}^{\top}\bm{W}_{H}\bm{H}_{D}}$
We tried this and ran into numerical problems, since the iterates might take on negative entries! To preserve nonnegativity, we need to be more careful and apply the approach in algorithm 1, derived from the work [28].
As a limitation, the paper [28] only outlined a justification of [14, Thm. 1] following the original proof strategy, but they never proved the method’s validity in detail. As another issue, the proof strategy in [14, Thm. 1] assumes positive-valued entries, but the algorithm should remain valid as long as the denominator terms are positive. To fill in these gaps, appendix A establishes a full theoretical guarantee for the MUR method that uses a slightly different argument from [14, Thm. 1] based on evaluating the best multiplicative (rather than additive) perturbations. This argument leads to the following result.
Theorem 2 (Convergence of MUR algorithm for partially fixed NMF).